Casey and Clarissa want to surprise their father and build a fence for the pig pen in the shape of a Triangle. They have 74' of fencing.
The second side of the fence is two times longer than the first side. The third side of the fence is 10 feet longer than the first side. What are the dimensions of the tringular fence.
I tried: x + 2x + 10 + x = 74
4x - 10 = 74 - 10
4x = 64
This is where I get lost, as to what calculatons come next. Do I divide?...the final numbers do not total 74! Help!

Answer by checkley71(8403)

good start!!!!!
4x=64 now divide both sides by 4
4x/4=64/4
x=16 for the short side.
2*16=32
16+10=26
proof
16=32+26=74
74=74